OpenStudy (anonymous):

use sigma notation to write the sum: (3/1+1)+(3/1+2)+(3/1+3)...+(3/1+14).

OpenStudy (anonymous):

i am going to guess that this is \[\frac{3}{1+1}+\frac{3}{1+2}+\frac{3}{1+3}+...\frac{3}{1+14}\]

OpenStudy (anonymous):

in which case you could us \[\sum_{k=1}^{14}\frac{3}{1+k}\] for example

OpenStudy (anonymous):

or even \[3\sum_{k=1}^{14}\frac{1}{1+k}\]

OpenStudy (anonymous):

yes! Thank you !!! on my choices i have this as an option...

OpenStudy (anonymous):

\[\sum_{j-1}^{14} 3\div1+j\]

OpenStudy (anonymous):

3/(1+j)

OpenStudy (anonymous):

would that still be correct?